# Counterfactuals As a Means for Evaluating Faithfulness of Attribution Methods in Autoregressive Language Models

### v3

@@ 2 Related work
Evaluating Explanations. Most current metrics for evaluating faithfulness involve either removing important tokens or retraining the model using only those identified as important by attribution methods (Chan et al., 2022). For instance, Abnar and Zuidema (2020) assess explanations by comparing them with gradient and ablation techniques. Although Wiegreffe and Pinter (2019) caution that gradients should not be considered ideal or the “ground truth,” they still utilize gradients as a proxy for the model’s intrinsic semantics. Importantly, the trustworthiness of explanations is both task- and model-dependent (Bastings et al., 2022), and different attribution methods frequently produce inconsistent results (Neely et al., 2022). As a result, it is not justifiable to treat any single explanation method as a universal standard across all contexts.
@@
Further, Han et al. (2020) and Jain et al. (2020) frame attribution methods as either faithful or unfaithful, with no consideration for degrees of faithfulness. They describe attribution methods that are "faithful by construction." In contrast, other researchers propose that faithfulness exists on a spectrum and suggest evaluating the “degree of faithfulness” of explanation methods (Jacovi and Goldberg, 2020). Our approach aligns with this view, as we aim to find explanation methods that are sufficiently faithful for autoregressive models.
Atanasova et al. (2023) evaluate the faithfulness of natural language explanations using counterfactuals, applying techniques from Ross et al. (2021) to assess how well explanations align with the model’s decision-making. This line of work offers valuable insights into the use of counterfactuals, which we build upon for evaluating attribution methods in language models. Another relevant direction is the evaluation of explanations using uncertainty estimation. For example, Slack et al. (2021) develop a Bayesian framework that generates feature importance estimates along with their associated uncertainty, expressed through credible intervals, highlighting the importance of uncertainty in faithfulness evaluations
The OOD Problem in Explainability.
The issue of OOD inputs in explainability has been raised by several works. Hooker et al. (2019) and Vafa et al. (2021) suggest retraining or fine-tuning the model using partially erased inputs to align training and evaluation distributions. However, this process can be computationally expensive and is not always practical. An alternative approach by Kim et al. (2020) aims to ensure that the explanation remains in-distribution to mitigate OOD problems. Our work addresses this concern by preserving the input distribution during faithfulness evaluation, particularly for autoregressive models.
@@ 5.1 The Out-of-Distribution Problem
The bottom matrix in Figure 4 illustrates the difference between the first and second matrices. It shows that the correlation difference between fine-tuned and instruct-tuned predictors is negligible when editors are used for token replacement. However, significant differences arise when other methods (e.g., <unk>/Erase/<mask>/att-zero) are employed, underscoring the importance of avoiding OOD inputs when evaluating explanations for off-the-shelf models.

### v4

@@ 2 Related work
Evaluating Explanations. One line of work uses another method as a proxy for evaluating an explanation. For instance, Abnar and Zuidema (2020) assess explanations by comparing them with gradient. Although Wiegreffe and Pinter (2019) caution that gradients should not be considered ideal or as the “ground truth,” they nonetheless utilize gradients as a proxy for the model’s intrinsic semantics.
Most current metrics for evaluating faithfulness involve either removing important tokens or retraining the model using only those identified as important by attribution methods (Chan et al., 2022). Importantly, the trustworthiness of explanations is both task- and model-dependent (Bastings et al., 2022), and different attribution methods frequently produce inconsistent results (Neely et al., 2022). As a result, it is not justifiable to treat any single explanation method as a universal standard across all contexts.
@@
Further, Han et al. (2020) and Jain et al. (2020) frame attribution methods as either faithful or unfaithful, with no consideration for degrees of faithfulness. They describe attribution methods that are “faithful by construction.” In contrast, other researchers propose that faithfulness exists on a spectrum and suggest evaluating the “degree of faithfulness” of explanation methods (Jacovi and Goldberg, 2020). Our approach aligns with this view, as we aim to find explanation methods that are sufficiently faithful for autoregressive models.
Atanasova et al. (2023) evaluate the faithfulness of natural language explanations using counterfactuals, applying techniques from Ross et al. (2021) to assess how well explanations align with the model’s decision-making. This line of work offers valuable insights into the use of counterfactuals, which we build upon for evaluating attribution methods in language models.
The OOD Problem in Explainability. The issue of OOD inputs in explainability has been raised by several works. Hooker et al. (2019) and Vafa et al. (2021) suggest retraining or fine-tuning the model using partially erased inputs to align training and evaluation distributions. However, this process can be computationally expensive and is not always practical. An alternative approach by Kim et al. (2020) aims to ensure that the explanation remains in-distribution to mitigate OOD problems. Our work addresses this concern by preserving the input distribution during faithfulness evaluation, particularly for autoregressive models.
@@ 5.1 The Out-of-Distribution Problem
The bottom matrix in Figure 4 illustrates the difference between the first and second matrices. It shows that the correlation difference between fine-tuned and instruct-tuned predictors is negligible when editors are used for token replacement. However, significant differences arise when other methods (i.e., <unk>/Erase/<mask>/att-zero) are employed, underscoring the importance of avoiding OOD inputs when evaluating explanations for off-the-shelf models.